OpenStudy (anonymous):

Explain? What is the value of n in the equation 32 × 33 = 3 n ? A. 3 B. 5 C. 6 D. 8

OpenStudy (kohai):

Oh, okay. So, first you're going to multiply the 32*33, and then divide by 3 to get n

OpenStudy (kohai):

So n=?

OpenStudy (anonymous):

there is something really wrong with this question

OpenStudy (anonymous):

agreed^^ did you mean 3^n?

OpenStudy (anonymous):

that is, none of your answer choices are remotely correct can you take a screenshot? i think there is a typo somewhere

OpenStudy (kohai):

There is, yeah

OpenStudy (anonymous):

you wrote \[32\times 33=3n\] which, when dividing by \(3\) would give you \[n=32\times 11\]

OpenStudy (anonymous):

oooh i bet it is this \[\large 3^2\times 3^3=3^n\] am i right?

OpenStudy (anonymous):

they are only asking you to add up the exponents to find \(n\)

OpenStudy (anonymous):

3^2 = 9 3^3 = 27 3^n =? I think your right @satellite73

OpenStudy (anonymous):

@satellite73 @mathbrain99 @aric200 @iPwnBunnies @satellite73 @mathlover2014 @mathstudent55 @girlnotonfire @johnweldon1993 @Hero @OrangeMaster @RatzLover95 @Squirrels @Whitemonsterbunny17 @undeadknight26 @whpalmer4 @Yuba

OpenStudy (anonymous):

Can someone please help? I have to leave in like 30 minutes to pick my mom up from work and I still have 8 more questions. :) I wanna know this stuff for High School. :)